OpenStudy (anonymous):

Prove that the given rule is a linear transformation, or show why it isn’t. (a) T : \[\mathbb{R}^\infty \rightarrow \mathbb{R}^\infty\] T is the rule “shift to the right”: T(x1, x2, x3, . . .) = (0, x1, x2, x3, . . .). (b) T : M2×3 \[(\mathbb{R}) \rightarrow \mathbb{R}^2\] T sends the matrix M to Mw where ~w = (1, 2, 3) (the output is a vector in \[\mathbb{R}^2\]. (c) T : \[W _{2} \rightarrow \mathbb{R}^2\], T(x, y) = (x, y). (d) T : \[W _{2} \rightarrow \mathbb{R}^2\], T(x, y) = (ln(x), ln(y)).

OpenStudy (anonymous):

(e) T : \[C^\infty (\mathbb{R}) \rightarrow \mathbb{R}, T(f) = \int\limits_{1}^{3} f \times \sin (x)dx\]